OpenStudy (anonymous):

Evaluate each expression. 3 x 4 ^ -4 over 3^-4 x 4 ^-1

OpenStudy (anonymous):

\[\frac{ \frac{ 3 }{ 4^{4} } }{ \frac{ 1 }{ 3^{4}4^{1} } }\] ??

OpenStudy (anonymous):